Systems Staff Meeting Notes
July 16, 2007
July 16, 2007
Present: Thomas Bielicke, Cynthia Bowker, Will Chaney, Logan Clapp, Winston Harris, Bobbie Meng, Tom Minton, Amy Polk, Mark Sullivan, Stephen WilliamsAnnouncements
Ongoing Systems Updates
- Per Bill, in today's Gainesville Sun, there is an article about Alachua County giving out prescription cards with a 30% discount.
- Meeting about the VOIP phones at 1PM today. There is another one scheduled for next Monday as well. There are three styles, each with individual numbers. All business calls should be routed to a main number.
- Orders that are $1000 or less will be put on hold, unless determined to be a necessary purchase.
- DLC - FTP project - GB of data -MS - two logical drives reserved for this: old Apaches. They are small drives. We will hold two as spares - either Linux or AV server.
- Tom's review: the movie "Transformers" was very good - and a long ad for Chevy Camaro.
- The new Harry Potter movie is not for kids.
- No Vista implementation until ALEPH v.18 is up and running.
- UF VPN client doesn't work in Vista. The trick is to right click and install as administrator, and then run as administrator.
- Public computer replacements - Evan reports that four out of seven have bad hard drives. Issuing machines - Documents moved CPUs to MSL. Will move them to Evan or MSL, then reissue to branch liaisons.
Software Projects Status
- The AV server is up - no complaints - client upgrade.
- Xerox - West - problems last week: printers not installed. There was no printer installed on half of the machines.
- Exchange synching with mobile devices - deleted all email and then it worked.
- SAN all racked - request for VM Ware training submitted.
- Vista upgrade for Systems - only Logan has an image. It is general; will create other parts later. The image has Office.
- SMS - going along.maybe this week.
- Server updates - Logan checked last week - they were fine.
- Saturday - the Exchange server failed over to Exchange 2 - then NT2R also failed over. NT2r1 was stuck - Will had to come in to fix it.
- Campus AD - no news.
- FYI.Logan had four terabyte drives from Western Digital - one was dead - originals - 2 SATA in RAID, this one was a power issue.
- Grover - Stephen and Chris are working on a design review and will present the design to staff prior to coding.
- Tracking Database - Test versions released to DLC for testing.
- Greenstone Production - died on Friday - software only. Finally decided it was the Apache updates. Backed out, reinstalled Greenstone software - and that seemed to fix it. It almost looked like we had been attacked - the box is very visible.
- Patron Updates - OPAC changes - sending v.18 files - including fall students 27,000 to 65,000.
- Software change to ITS - local change to show when classes are full.
- NLBP database - Lori requested an application to list all Library barcode prefixes. Should be quick. National interest. Stephen found a few lists out there - OCLC symbol list exists - filler task.
Student & Department
- Endeca is a bust - statewide test crashed it. It is used in large corporations and NC State did good things with it.
- ALEPH v.18 interface - trying to reach an agreement. Carol Turner likes the look of UNF's.
- Satisfaction survey posted this afternoon - look on the homepage.
- New UF template has one orange line only.
- Trudi is retiring at the end of the year - phased retirement. The deal is that her line can't be filled.
Systems Security Management
- New student, Mindy Stephenson, has started working.
- Meeting held.